Determine the density of sulphur dioxide , SO2 gas in g/L at 25 deg C and 100 kPa 

1
Expert's answer
2021-06-23T05:20:25-0400

Molar Mass of SO2 = 64.066

1 kPa = 0.00987atm

= 100× 0.00987

= 0.987atm

T = 273+25 = 298K

Density = (64.066 × 0.987)/(0.0821×298)

= 63.233/24.466

= 2.58g/L